OpenStudy (anonymous):

What does ¹3/4+1/2+¹1/4=?

OpenStudy (anonymous):

What are the subscript 1's? Are you trying to say \[1 \frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

Or is your question just 3/4+1/2+1/4

OpenStudy (anonymous):

Yes it means the first comment you put

OpenStudy (anonymous):

Oh well \[1 \frac{ 3 }{ 4 } \implies 1+\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

Do you know how to add that or do you need help with fractions in general?

OpenStudy (anonymous):

I just need hep with this particular question I know has to add and all that other stuff I'm just stuck on this problem for some reason

OpenStudy (anonymous):

Oh ok, can you show me specifically what place you're stuck at? Then I can see where you went wrong.

OpenStudy (anonymous):

Oh um...never mind I get now my big bro just showed how to do it!! But thanks anyways

OpenStudy (anonymous):

Np, glad you understood it :), take care.